OpenStudy (anonymous):

for medal: how do you find the derivative of g(x)= |2x+3| those are absolute value signs.

OpenStudy (raffle_snaffle):

power rule

OpenStudy (anonymous):

is the answer just 2?

OpenStudy (raffle_snaffle):

yeah

OpenStudy (anonymous):

do the absolute value signs affect how you take the derivative?

OpenStudy (raffle_snaffle):

why would they?

OpenStudy (anonymous):

i don't know

OpenStudy (raffle_snaffle):

do you know what abs signs mean?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what do they mean?

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

satellite, the answer is not 2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

\[f(x) = |2x+3| = \left\{\begin{array}{rcc} -2x-3& \text{if} & x < -\frac{3}{2} \\ 2x+3& \text{if} & x \geq -\frac{3}{2} \end{array} \right. \]

OpenStudy (anonymous):

so your derivative is \[f'(x) = \left\{\begin{array}{rcc} -2& \text{if} & x < -\frac{3}{2} \\ 2& \text{if} & x > -\frac{3}{2} \end{array} \right.\]

OpenStudy (anonymous):

and also it is not defined at \(x=-\frac{3}{2}\) since the limit does not exist there, i.e. since \(-2\neq 2\)

OpenStudy (anonymous):

absolute value is a piecwise function, that is why, even though it seems easy when you first see it in grade school, it is a real pain in the arse later on

OpenStudy (anonymous):

i'm looking at my textbook and it says that a critical number exists when the derivative=0 or when the derivative does not exist (like u said earlier).

OpenStudy (anonymous):

thank you. it makes sense now.